# EEG Emotion Copilot: Optimizing Lightweight LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

### v2

@@ III-D Model Deployment in EEG Emotion Copilot
Algorithm 2 Pruning Process
1:Input: Initial model $\mathcal{M}$, Parameter group set $\mathcal{P^{\prime}}$, importance function $\mathcal{I}$, contraction strength $\alpha$
2:Output: Pruned model $\mathcal{M}^{\prime}$
3:Calculate group importance:
4:for all $p^{\prime}\in\mathcal{P^{\prime}}$ do
5:     $\mathcal{I}(p^{\prime})\leftarrow\sum_{w\in\mathcal{P^{\prime}}}\|w\|_{2}^{2}$ (See Formula 3, 4, 5, 6)
6:end for
7:    $\mathcal{P}^{\prime\prime}\leftarrow\text{sort}(\mathcal{P^{\prime}},\text{by }\mathcal{I})$
8:Select groups for pruning based on importance:
9:for all $p^{\prime\prime}\in\mathcal{P}^{\prime\prime}$ do
10:     if $\mathcal{I}(p^{\prime\prime})$leq $\theta$ is satisfied then #$\theta$ is the threshold
11:         $\mathcal{P^{\prime\prime}}\leftarrow\mathcal{P^{\prime\prime}}\setminus p^{\prime\prime}$
12:     end if
13:end for
14:return $\mathcal{M}^{\prime}$
III-E Generation of Assisted Electronic Medical Records
Due to the fact that the proposed copilot involves various tasks such as emotion recognition, diagnosis, and treatment plan design, it is sufficient to embed demographic data from different cases under various emotional states, along with the corresponding compressed EEG signals, into the training data. In downstream applications, participant data can be input into the fine-tuned model in a specific format to obtain diagnostic information and treatment plans, as shown in Fig. 7.
@@ III-E Generation of Assisted Electronic Medical Records
[…]
IV EXPERIMENTS
IV-A Implementation Details
Dataset: We utilize the FACED dataset [29] to validate our copilot. This dataset includes recordings from 123 subjects, each with 32 EEG channels, as they watched 28 video clips, each lasting 30 seconds. These clips encompass nine emotions: anger, disgust, fear, sadness, neutral, amusement, Inspiration, Joy, and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 and then trained on the public datasets with a learning rate of 1e-5, incorporating L1 and L2 regularization terms. The training was conducted for 40 Epoches. During the LoRA fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRAr to 8, LoRAα to 32, and LoRAdropout to 0.5. The fine-tuning process spanned 15 epochs, dividing into 5 steps. Model pruning was executed on a NVIDIA A800 80G, while the fine-tuning were completed on NVIDIA GeForce RTX 4090 GPUs.
IV-B Performance Validation of Varied EEG Signal Length
@@
TABLE I: Performance Comparison Of Competing Light-Weight Models
@@
base_model-2 [22] 0.50B 0.920G bf16 6 0.153 1.230 0.450 1.290
@@
[…]
IV-C Performance Comparison with Lightweight Models
@@
To better validate our experimental hypothesis, we combine the strategies outlined in Fig. 6 and Fig. 8. First, we divide the training process based on the data derived from compression method 1, into five steps, with each step trained for 3 epochs. For each step’s checkpoint (i.e., checkpoint1-n), we perform classification of content generation for the nine emotions described in III-A, followed by classification for three emotions (positive, negative, neutral). For comparison, we use Qwen2-0.5B as the base model, training it for 6 epochs under identical experimental conditions to obtain base_model-1, and perform classification, as shown in Table I.
@@
Next, we fine-tuned the models based on the data derived from compression method 2. Specifically, we fine-tuned the pruned model in three steps, with 3, 3, and 3 epochs respectively, and perform classification for both the nine emotions and the three emotions. For consistency, Qwen2-0.5B is also trained for 6 epochs to obtain base_model-2, after which the corresponding classification experiment is conducted, as shown in Table I.
@@
Additionally, we include opt-350m [58], LiteLlama, and gpt-sw3-356m as baseline references. Indeed, the checkpoint1-n model of Model 1 did not achieve ideal results on both tasks compared to the base_model-1. However, considering the F1 scores across multiple epochs of training, the model’s performance on this task gradually improved. In Model 2, the performance of checkpoint2-2 and checkpoint2-3 showed significant improvement compared to models with the similar parameter size, and they also consumed less average response time.
IV-D Performance Comparison with Heavyweight Models
@@
In Table II, we present the results of emotion recognition for Qwen2-1.5B [22], InternLM2.5-1.8B [59], Qwen2.5-3B [60], and Qwen2.5-7B [60], and InternLM2.5-7B [59]. Surprisingly, the results were not as impressive as expected. The reason for this is that the default setting for top-k is 50. But even if top-k is set to 1, the results are still not ideal. For the heavyweight models, their responses are more flexible due to their larger knowledge storage, allowing them to generate a wider range of answers. This also indicates that, despite the very low loss during training, which might suggest model fitting, the actual responses in specific scenarios are dissatisfactory.
Figure 10: Clustering results of data for compression method 1 and 2 described in Fig. 8 using various models mentioned in Table I and II with K-Means clustering and t-SNE [61] visualization of the text embeddings.
Furthermore, we first use the pre-trained all-MiniLM-L6-v2 model for text representation. Specifically, we combine the “prompt” and “treatment” from each record into a single string and generate the corresponding embedding vectors. These embeddings are then subjected to K-Means clustering and spectral clustering [62] for unsupervised learning, with the number of clusters set to 9. The t-SNE [61] is applied to reduce the dimensionality of the embedding and visualize the results, as shown in Fig. 9.
In addition, we investigate the clustering performance for data of compression method 1 and 2 respectively in Fig. 8 by applying the same procedure to a range of models included in Table I and II. We convert the text into the appropriate input format using the tokenizer, extract the hidden states from the final layer, and apply average pooling to generate fixed-length text embedding vectors. These embeddings are clustered using K-Means, and dimensionality reduction is performed using the t-SNE for visualization, as shown in Fig. 10. Indeed, as illustrated in Fig. 9 and 10, the visualization suggests that corpora of two different lengths are not easily clustered. Specifically, compared to the data of Model 1, the data of Model 2 exhibits a higher degree of aggregation, which directly corresponds to a reduction in computational load.
IV-E Model Fine-tuning Strategy Comparison After Pruning
To evaluate the two fine-tuning strategies outlined in Fig. 5, we adopted compression method 2 show in Fig. 8 for comparison. According to the results from Table I and Fig. 11, we found that Strategy 1 demonstrated better performance in Table I than Strategy 2 did in the lightweight models fine-tuning process.
TABLE II: Performance comparison Of Heavy-Weight Models
@@ IV-E Model Fine-tuning Strategy Comparison After Pruning
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 5.870 0.035 5.910
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.111 4.260 0.041 10.960
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.051 1.410 0.152 2.100
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.082 10.090 0.039 10.090
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.074 3.730 0.328 3.510
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 6.320 0.041 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.133 3.680 0.114 3.760
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.036 0.020 0.038 0.240
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.082 10.260 0.042 7.300
@@
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 7.100 0.036 5.830
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.080 3.810 0.342 4.500
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.040 0.690 0.036 0.870
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.090 8.950 0.040 9.390
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.081 3.780 0.292 4.170
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 7.210 0.036 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.036 6.320 0.038 4.470
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.051 4.870 0.110 4.920
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.081 11.640 0.041 9.790
IV-F Assisted Electronic Medical Record Generation with Our Copilot
Fig. 12 illustrates an example of using our EEG Emotion Copilot to create an electronic medical record. Initially, the starting sentence of the prompt is provided, followed by the addition of demographic information of the subjects, including age, gender, and certain facial features. Finally, the preprocessed and compressed signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot. The EEG Emotion Copilot first analyzes the data to generate an answer, which is subsequently transformed into the record through format conversion.
V Discussion
V-A Data Structure of Prompt

### v3

@@ 3.5 Generation of Assisted Electronic Medical Records
4 EXPERIMENTS
4.1 Implementation Details
Dataset: During pruning, the Wikitext dataset is used to optimize language capabilities. For training validation, we utilize the FACED dataset [30], which contains EEG recordings from 123 subjects (75 females, mean age 23.2) with 32 channels, collected via a NeuSen.W32 system at 250 or 1000 Hz. Participants watched 28 affective video clips (30 seconds each) designed to evoke nine emotions: anger, disgust, fear, sadness, neutral, amusement, inspiration, joy, and tenderness. Each positive and negative emotion is represented by three clips, and neutral by four, ensuring balanced coverage.
We further adopted the SEED [31, 28] and SEED-IV [65] datasets, released by Shanghai Jiao Tong University, to validate the proposed method. Both datasets contain EEG recordings from 15 healthy subjects across three sessions on different days, acquired using a 62-channel ESI NeuroScan system (originally sampled at 1000 Hz and downsampled to 200 Hz). In SEED, participants viewed 15 film clips designed to elicit positive, neutral, or negative emotions. SEED-IV extends the emotion categories to happiness, sadness, fear, and neutral, and additionally provides eye-tracking data recorded via SMI eye-tracking glasses. Each trial includes a 5-second cue,  4-minute clip, 45-second self-assessment, and 15-second rest.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 with Wikitext dataset used to optimize language capabilities and then trained on the public datasets with a learning rate of 1e-5, incorporating L1 and L2 regularization terms.The training was conducted for 40 Epoches. During the LoRA fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRAr to 8, LoRAα to 32, and LoRAdropout to 0.5. The fine-tuning process spanned 15 epochs, dividing into 5 steps. The training epochs of the compared models were set to 15. Unlike prior work requiring high-end GPUs (e.g., RTX 4090) for deployment, our 0.5B model runs on standard hardware, though training multiple models (0.5B to 7B) necessitated A800 and RTX 4090 GPUs for memory-intensive computations. Besides, all other models were trained for 15 epochs to ensure fairness and comparability.
4.2 Performance Validation of Varied EEG Signal Length
@@
4.3 Training details of different pruning ratio
We investigated the parameter distributions of the pruned models under various pruning ratios, as illustrated in Table 1 and Fig. 8. Pruning effectively reduces the dimensionality of key components, including the MLP layers (gate_proj, up_proj, down_proj) and the attention modules (q_proj, k_proj, v_proj, o_proj), resulting in smaller matrix operations during both forward and backward passes. These structural changes significantly reduce the number of floating-point operations and inference time.
In addition, we explored high pruning ratios of 0.9, 0.8, 0.7, and 0.6 for the Qwen2-0.5B model. However, training under pruning ratios of 0.9 and 0.8 consistently failed to converge, as shown in Fig. 9, likely due to excessive removal of critical parameters. While models with pruning ratios of 0.7 and 0.6 exhibited more stable convergence, their improvements remained marginal after repeated training attempts. We attribute this to the relatively higher proportion of retained parameters, which preserved a basic level of representational capacity. Nevertheless, the aggressive pruning still disrupted essential pathways, limiting the model’s ability to recover, as reflected in the flattened learning curves.
Figure 8: Module parameter trends across different pruning ratios.
Figure 9: The training loss of different pruning ratio (0.9, 0.8, 0.7, 0.6).
Table 1: Changes in Model Parameter Dimensions at Different Pruning Ratio
Pruning Ratio Embed Dim Q Proj K/V Proj MLP Gate/Up MLP Down Layers Norm Eps Total Params GPU memory usage for loading (MB) GPU memory usage while pruning (MB)
Full 896 896 128 4864 896 24 1e-6 498,431,872 1628 -
0.1 804 804 134 4376 804 24 1e-6 411,738,852 2546 2837
0.2 716 710 142 3888 716 24 1e-6 338,560,828 1406 1731
0.3 624 624 156 3404 624 24 1e-6 271,158,576 1242 2543
0.4 536 536 134 2916 536 24 1e-6 211,255,288 1090 2271
0.5 448 444 148 2432 448 24 1e-6 159,284,000 868 2451
0.6 356 356 178 1944 356 24 1e-6 113,233,828 832 2903
0.7 268 268 134 1456 268 24 1e-6 74,699,660 866 1731
0.8 176 168 24 972 176 24 1e-6 41,563,888 704 2291
0.9 88 84 12 484 88 24 1e-6 17,283,320 586 2273
Table 2: Performance Comparison Of Competing Light-Weight Models On FACED
@@ 4.3 Training details of different pruning ratio
checkpoint2-4 0.15B 0.294G bf16 3 0.346 0.860 0.990 0.550
checkpoint2-5 0.15B 0.294G bf16 3 0.351 0.900 0.991 0.570
base_model-2 [23] 0.50B 0.920G bf16 15 0.120 1.504 0.970 1.811
@@
4.4 Performance Comparison with Lightweight Models
@@
To rigorously validate our experimental hypothesis, we integrate the strategies illustrated in Fig. 6 and Fig. 10. The training process is structured based on data derived from compression method 1, divided into five sequential steps, with each step trained for three epochs. At each checkpoint (i.e., checkpoint1–n), we conduct classification on content generation for the nine emotion categories described in 3.1, followed by a broader classification into three sentiment groups (positive, negative, and neutral). For comparative analysis, we adopt Qwen2-0.5B as the baseline model, training it for six epochs under identical experimental conditions to obtain base_model-1, which is then evaluated for classification performance, as summarized in Table 2.
@@
Next, we fine-tuned the models based on the data derived from compression method 2. Specifically, we fine-tuned the pruned model in 5 steps, with 3 epochs respectively, and perform classification for both the nine emotions and the three emotions. For consistency, Qwen2-0.5B is also trained for 15 epochs to obtain base_model-2, after which the corresponding classification experiment is conducted, as shown in Table 2.
@@
Additionally, we include opt-350m [66], LiteLlama, and gpt-sw3-356m and recently released qwen3-0.6B [67] as baseline references. Indeed, the checkpoint1-n model of Model 1 did not achieve ideal results on both tasks compared to the base_model-1. However, considering the F1 scores across multiple epochs of training, the model’s performance on this task gradually improved. In Model 2, the performance of checkpoint2-2 and checkpoint2-3 showed significant improvement compared to models with the similar parameter size, and they also consumed less average response time.
4.5 Performance Comparison with Heavyweight Models
@@
In Table 3, we present the results of emotion recognition for Qwen2-1.5B [23], InternLM2.5-1.8B [68], Qwen2.5-3B [69], and Qwen2.5-7B [69], and Internlm2.5-7B [68] and DeepSeek-R1-Distill-Qwen-1.5B [70]. Surprisingly, the results were not as impressive as expected. The reason for this is that the default setting for top-k is 50. But even if top-k is set to 1, the results are still not ideal. For the heavyweight models, their responses are more flexible due to their larger knowledge storage, allowing them to generate a wider range of answers. This also indicates that, despite the very low loss during training, which might suggest model fitting, the actual responses in specific scenarios are dissatisfactory.
Table 3: Performance comparison Of Heavy-Weight Models On FACED
@@
Internlm2.5-1.8B-1 [68] 1.8B 3.780G bf16 15 0.111 4.26 0.931 4.95
qwen2.5-3b-1 [69] 3B 6.170G bf16 15 0.272 5.17 0.935 5.15
DS-R1-Dis-Qwen-1.5B-1 [70] 1.5B 3.55G bf16 15 0.143 15.28 0.950 14.87
qwen2.5-7b-1 [69] 7B 15.230G bf16 15 0.238 6.30 0.945 6.18
Internlm2.5-7B-1 [68] 7B 14.200G bf16 15 0.106 14.89 0.978 15.29
W qwen2-1.5B-2 [23] 1.5B 3.091G bf16 15 0.117 3.48 0.993 3.03
Internlm2.5-1.8B-2 [68] 1.8B 3.780G bf16 15 0.110 15.42 0.970 4.89
qwen2.5-3b-2 [69] 3B 6.170G bf16 15 0.201 10.24 0.932 5.15
DS-R1-Dis-Qwen-1.5B-2 [70] 1.5B 3.55G bf16 15 0.156 14.28 0.970 14.43
qwen2.5-7b-2 [69] 7B 15.230G bf16 15 0.179 5.05 0.990 6.10
Internlm2.5-7B-2 [68] 7B 14.200G bf16 15 0.169 14.89 0.962 14.92
Random-guess - - - - 0.111 - 0.333 -
@@
Internlm2.5-1.8B-1 [68] 1.8B 3.780G bf16 15 0.104 6.44 0.960 4.95
qwen2.5-3b-1 [69] 3B 6.170G bf16 15 0.153 5.34 0.822 5.28
DS-R1-Dis-Qwen-1.5B-1 [70] 1.5B 3.55G bf16 15 0.140 9.10 0.970 5.80
qwen2.5-7b-1 [69] 7B 15.230G bf16 15 0.226 6.24 0.960 6.14
Internlm2.5-7B-1 [68] 7B 14.200G bf16 15 0.106 14.88 0.967 15.25
Random-guess - - - - 0.111 - 0.333 -
W qwen2-1.5B-2 [23] 1.5B 3.091G bf16 15 0.093 3.45 0.994 3.13
Internlm2.5-1.8B-2 [68] 1.8B 3.780G bf16 15 0.112 6.09 0.956 4.90
qwen2.5-3b-2 [69] 3B 6.170G bf16 15 0.178 9.80 0.844 5.03
DS-R1-Dis-Qwen-1.5B-2 [70] 1.5B 3.55G bf16 15 0.146 8.48 0.989 8.34
qwen2.5-7b-2 [69] 7B 15.230G bf16 15 0.153 6.53 0.970 6.17
Internlm2.5-7B-2 [68] 7B 14.200G bf16 15 0.168 14.95 0.967 7.46
Random-guess - - - - 0.111 - 0.333 -
Note: DS-R1-Dis-Qwen-1.5B denotes DeepSeek-R1-Distill-Qwen-1.5B.
Table 4: Performance Comparison Of Competing Light-Weight Models and Heavy-Weight Models On SEED
Compression Model Parameters Storage_size Precision Epoch Three Emotion Recognition
Method F1 (Top-K=50) Avg. RT (s) F1 (Top-K=1) Avg. RT (s)
W Our Model
checkpoint1 0.15B 0.294G bf16 15 0.570 0.74 0.607 0.75
checkpoint2 0.15B 0.294G bf16 15 0.630 1.13 0.640 1.26
checkpoint3 0.15B 0.294G bf16 15 0.710 1.10 0.720 1.03
checkpoint4 0.15B 0.294G bf16 15 0.768 1.25 0.772 1.05
checkpoint5 0.15B 0.294G bf16 15 0.821 1.12 0.835 0.95
base_model [23] 0.50B 0.920G bf16 15 0.574 2.17 0.537 1.65
opt-350m [66] 0.35B 0.647G bf16 15 0.169 0.98 0.193 0.95
LiteLlama 0.46B 0.901G bf16 15 0.628 1.12 0.640 1.16
gpt-sw3-356m 0.47B 1.580G bf16 15 0.315 0.84 0.337 0.81
qwen3-0.6B [67] 0.6B 1.5G bf16 15 0.597 33.20 0.601 31.73
W qwen2-1.5B [23] 1.5B 3.091G bf16 15 0.533 2.25 0.521 2.12
Internlm2.5-1.8B [68] 1.8B 3.780G bf16 15 0.635 3.12 0.640 3.23
qwen2.5-3b [69] 3B 6.170G bf16 15 0.573 3.67 0.605 3.07
DS-R1-Dis-Qwen-1.5B [70] 0.6B 3.55G bf16 15 0.566 38.19 0.571 36.17
qwen2.5-7b [69] 7B 15.230G bf16 15 0.581 4.42 0.596 5.22
Internlm2.5-7B [68] 7B 14.200G bf16 15 0.633 11.46 0.637 11.63
Random-guess - - - - 0.333 - 0.333 -
[…]
Table 5: Performance Comparison Of Competing Light-Weight Models and Heavy-Weight Models On SEED-IV
Compression Model Parameters Storage_size Precision Epoch Three Emotion Recognition
Method F1 (Top-K=50) Avg. RT (s) F1 (Top-K=1) Avg. RT (s)
W Our Model
checkpoint1 0.15B 0.294G bf16 15 0.870 1.05 0.860 1.15
checkpoint2 0.15B 0.294G bf16 15 0.980 1.03 0.990 1.08
checkpoint3 0.15B 0.294G bf16 15 0.990 1.02 0.995 1.12
checkpoint4 0.15B 0.294G bf16 15 0.970 1.13 0.974 1.02
checkpoint5 0.15B 0.294G bf16 15 0.975 1.17 0.989 1.14
base_model [23] 0.50B 0.920G bf16 15 0.995 2.01 0.995 1.72
opt-350m [66] 0.35B 0.647G bf16 15 0.287 0.77 0.246 1.24
LiteLlama 0.46B 0.901G bf16 15 0.920 1.02 0.921 1.14
gpt-sw3-356m 0.47B 1.580G bf16 15 0.819 0.87 0.755 0.85
qwen3-0.6B [67] 0.6B 1.5G bf16 15 0.269 60.26 0.273 69.83
W qwen2-1.5B [23] 1.5B 3.091G bf16 15 0.972 2.17 0.990 2.07
Internlm2.5-1.8B [68] 1.8B 3.780G bf16 15 0.940 6.96 0.945 7.09
qwen2.5-3b [69] 3B 6.170G bf16 15 0.981 3.56 0.991 3.15
DS-R1-Dis-Qwen-1.5B [70] 0.6B 3.55G bf16 15 0.454 66.35 0.437 45.86
qwen2.5-7B [69] 7B 15.230G bf16 15 0.986 4.85 0.995 4.83
Internlm2.5-7B [68] 7B 14.200G bf16 15 0.972 12.50 0.976 12.63
Random-guess - - - - 0.250 - 0.250 -
[…]
Figure 12: Clustering results of data for compression method 1 and 2 described in Fig. 10 using various models mentioned in Table 2 and 3 with K-Means clustering and t-SNE [71] visualization of the text embeddings.
Meanwhile, we also conducted experiments on the SEED and SEED-IV datasets using Compression Method 2. As shown in Table 4 and Table 5, this method achieves consistent performance gains under both intra-subject and cross-subject settings. Compared with unpruned models and other baselines, the proposed pruning strategy effectively reduces model size and inference latency while preserving or even improving classification accuracy on emotion recognition tasks.
Furthermore, we use the pre-trained all-MiniLM-L6-v2 model for text representation on FACED. Specifically, we combine the “prompt” and “treatment” from each record into a single string and generate the corresponding embedding vectors. These embeddings are then subjected to K-Means clustering and spectral clustering [72] for unsupervised learning, with the number of clusters set to 9. The t-SNE [71] is applied to reduce the dimensionality of the embedding and visualize the results, as shown in Fig. 11.
In addition, we investigate the clustering performance for data of compression method 1 and 2 respectively in Fig. 10 by applying the same procedure to a range of models included in Table 2 and 3. We convert the text into the appropriate input format using the tokenizer, extract the hidden states from the final layer, and apply average pooling to generate fixed-length text embedding vectors. These embeddings are clustered using K-Means, and dimensionality reduction is performed using the t-SNE for visualization, as shown in Fig. 12. Indeed, as illustrated in Fig. 11 and 12, the visualization suggests that corpora of two different lengths are not easily clustered. Specifically, compared to the data of Model 1, the data of Model 2 exhibits a higher degree of aggregation, which directly corresponds to a reduction in computational load.
4.6 Model Fine-tuning Strategy Comparison After Pruning
To rigorously assess the effectiveness of the two fine-tuning strategies illustrated in Fig. 5, we employed compression method 2, as depicted in Fig. 10, as a basis for comparison. As evidenced by the results in Table 2 and Fig. 13, Strategy 1 exhibited superior performance over Strategy 2 in Table 2, particularly in the fine-tuning process of lightweight models, underscoring its efficacy in optimizing model adaptation. Besides, Additionally, we visualize the training time and GPU memory usage of various models under the two compression methods presented in Tables 2 and 3, as illustrated in Fig. 14. Building upon Compression Method 2, and as illustrated in Tables 2 and 3, we further evaluated the feasibility of deploying various models on edge devices. While Tables 2 and 3 report the average runtime (Avg. RT) on an NVIDIA A800 GPU, practical deployment scenarios often involve resource-constrained environments where inference must be performed on CPUs alone. To this end, we deployed all models on a standard consumer-grade CPU (Intel i5-11400, 16GB RAM). The corresponding inference time and memory usage results are summarized in Table 6.
[…]
4.7 Assisted Electronic Medical Record Generation with Our Copilot
Fig. 15 presents a representative example of how our EEG Emotion Copilot can assist in generating electronic medical records. The process begins with an initial prompt containing a starting sentence, followed by the integration of demographic information such as the subject’s age, gender, and salient facial features. Subsequently, the preprocessed and compressed EEG signals are appended to complete the prompt, which is then fed into the EEG Emotion Copilot. The model first performs multimodal analysis to generate an appropriate response, which is subsequently structured into a standardized medical record through format conversion. To further enhance clinical decision support, we leverage recent advances in inference models such as DeepSeek and Qwen3 [73, 74, 75], which have demonstrated strong performance across several reasoning tasks. Building upon this, we endeavor to explore the performance of lightweight models in this context. In particular, we apply structured pruning (with a pruning ratio of 0.5) to the Qwen3-0.6B model to reduce complexity while preserving inference quality. We conduct extensive training for 25 epochs on the mental_health_counseling_conversations_sharegpt dataset-MHCCS, with 20% of the data for evaluation. Comparative experiments involving Qwen3-0.6B (pruned and unpruned), Qwen3-1.7B, and Qwen3-4B are conducted, with performance results illustrated in Fig. 16.
5 Discussion and Future Work
5.1 Data Structure of Prompt